OpenStudy (anonymous):

Write f(x) in standard form by completing the square. f(x)= -5x^2 + 80x - 318

OpenStudy (anonymous):

???

OpenStudy (anonymous):

X=8+or-(sqroot10)/5

OpenStudy (jim766):

-5x^2+ 80x - 318 = 0 1st get the 318 out of there +318 +318 -5x^2+80x = 318 now find Gcf of 5x^2 + 80x -5(x^2-16x) = 318 now complete the square

OpenStudy (jim766):

-5(x^2-16x) = 318 now complete the square -16/2 = -8 , now square it = 64 -5(x^2-16x + 64) = 318 + 64 -5(x^2-8)^2= 382

OpenStudy (jim766):

now move the 382 back over f(x) = -5(x-8)^2 - 382 does this make sense?

OpenStudy (anonymous):

no not really what would the answer be?

OpenStudy (jim766):

that is the answer it is in vertex form

OpenStudy (anonymous):

My answer is in standard form

OpenStudy (jim766):

yeah same thing

OpenStudy (anonymous):

thanks alot guys